Fermi National Laboratory


Accelerator Update

Monday June 14
The day shift began with Operations monitoring store 3571, with the Antiproton Source (Pbar) conducting studies, and with MiniBooNE and Switchyard (SY) taking beam.

Operations terminated store 3571 at 9:02 AM.

Operations prepared the Collider Detector at Fermilab (CDF) for a one-hour access beginning at 9:25 AM.

Operations put the Tevatron (TeV) through a dry squeeze at 11:02 AM.

At 11:21 AM, Operations reported that Booster was losing beam during injection. Operators investigated.

Operations put the TeV into shot setup at 12:32 PM. All SY beam was disabled.

At 1:27 PM, a Pbar Debuncher quadrupole power supply tripped off on an overcurrent indication. Experts investigated.

Operations established store 3574 at 3:17 PM with an initial luminosity of 78.33E30. This store set a NEW LUMINOSITY RECORD. Pbar experts conducted a twenty-minute aperture measurement study before returning to stacking.

Operations resumed sending beam to the Meson MCenter experiment.

A TeV RF station (TRF3) tripped off at 4:15 PM.

A Pbar Accumulator RF station (ARF1) tripped off at 4:55 PM due to an arcing problem. The Run Coordinator approved accessing the tunnel to repair the station. Operators prepared the area for access and technicians went in at 6:08 PM.

Operations began sending beam to the Meson MTest experiment at 6:44 PM. They finished using beam for the night at 7:48 PM.

The RF techs completed their repairs on ARF1 at 8:16 PM. Stacking resumed at 9:05 PM.

Earlier, CDF complained about the colliding beam problems. Operations reported at 9:23 PM that they had made collimator moves and talked with experts, but couldn't seem to make any improvements. Operators thought that a bad readback might be part of the problem.

At 9:31 PM, operators adjusted the vertical tune, and the horizontal and vertical Chromaticities. They moved collimators into the beam until monitors showed the effects.

At 10:52 PM, operators made more collimator adjustments, but they did not help.

Tuesday June 15
The midnight shift began with Operations monitoring store 3574, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 12:57 AM, Operations put the Recycler in to shot setup to stash antiprotons from Pbar.

At 1:14 AM, operators reported that they were making single collimator moves. At 1:48 AM, operators reported that this method showed a little success, but not enough to allow CDF to turn back on.

Recycler completed their stash and Pbar resumed stacking at 2:16 AM.

Operations reestablished beam to MCenter at 2:30 AM.

At 4:06 AM, a CDF expert expressed concern about not really knowing how much beam was actually in the beam abort gap, which was the cause of the colliding beam problem.

At 5 AM, Operations made further collimator and tune adjustments, but improved the colliding beam for CDF only slightly.

Operations talked with a TeV expert at 6:55 AM and tweaked the tunes, which reduced the beam abort gap problem a little.

At 8:30 AM, a TeV RF station (TRF5) jumped causing the abort gap counter to spike. An expert investigated and said there was a problem with the stations modulator, but since TRF3 was offline he couldn't do much until the store ended.

MCenter reported at 9:22 AM that they were finished with beam. Operations disabled all SY beam.

TeV experts began conducting an abort gap study at 10:45 AM by making small Tevatron Electron Lens (TEL) changes.

At 3:42 PM, the TeV suffered a sector F1 quench due to a CAMAC crate losing its -24 volt power supply. A tech replaced the supply.

CDF and D-zero (D0) accessed their collision halls at 5 PM.

TeV RF experts reported that they had completed the TeV RF repairs at 5:03 PM.

CDF completed their access at 5:21 PM and D0 finished their access at 6:13 PM.

Operations put the TeV through a dry squeeze at 6:50 PM and into shot setup at 8:08 PM.

A problem in Pbar's AP10 area at 9:21 PM put the TeV shot on hold. An operator had to reboot some equipment.

Operations established store 3576 at 11:04 PM with an initial luminosity of 71.52E30. CDF complained about a colliding beam proton problem. Operators consulted with TeV experts and adjusted the TeV collimators. This initially helped, but the problem returned. Operators again consulted with TeV experts.

Monday June 16
The midnight shift began with Operations monitoring store 3576 and the Recycler stash, with Operations trying to fix the colliding beam problems, with Pbar stacking, with MiniBooNE taking beam, and with operators establishing beam to MCenter.

At 12:38 AM, TeV experts came in to help operators resolve the colliding beam proton problem.

Operations established beam to MCenter at 1:04 AM.

With the TeV experts' help, CDF was able to turn on at 1:16 AM.

In preparing for a Recycler stash, Operations disabled SY beam at 2:26 AM.

The Recycler began stashing antiprotons at 3:41 AM. They finished and Pbar resumed stacking at 4:08 AM. Operators resumed sending beam to MCenter.

Plans for Today
MI experts will be allowed to conduct studies after the Pbar stack reaches 100 mA. TeV experts will conduct end-of-store studies late on the midnight shift and then Operations will terminate the store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/16/2004   email Fermilab